OpenStudy (anonymous):

< MEDAL > Kaitlyn can mow a lawn in 8 hours. Lance and Kaitlyn can mow the sam lawn together in 3 hours. How long would it take for Lance to mow the lawn by himself? a. 3h 23m b. 4h 48m c. 6h d. 7h

OpenStudy (anonymous):

1/8 + 1/x = 1/3 solve for x

OpenStudy (anonymous):

so 4 h & 48m ? @sourwing

OpenStudy (paki):

@ sourwing is rite... just make an equation like that, from the statement of the question and solve... 1/8 + 1/x = 1/3 1/x=1/3-1/8 take lcm and find "x"